OpenStudy (anonymous):

Solve the equation.6x over 7 minus x over 2 = 5

OpenStudy (anonymous):

\[\frac{6x}{7}-\frac{x}{2}=5\] \[\text{Taking LCM and adding } \frac{12x}{14}-\frac{7x}{14}=5\] Cross multiply 5x=70 x=14